Freight train A goes from point A to point B at 40 miles per hour. Freight train B takes 2 hours longer to go from point A to point B because it only goes 30 miles an hour. What is the distance from point A to point B?
Answer by stanbon(57239)

Freight train A goes from point A to point B at 40 miles per hour.
distance = x miles ; rate = 40 mph ; time = x/40 hrs.
------------------------------------------------------------
Freight train B takes 2 hours longer to go from point A to point B because it only goes 30 miles an hour.
distance = x miles ; rate = 30 mph ; time = x/30 hrs
---------------------------------------------
What is the distance from point A to point B?
Equation:
slower train - faster train = 2 hrs
x/30 - x/40 = 2 hrs
40x - 30x = 2*40*30
10x = 2*40*30
x = 240 miles
====================
Cheers,
Stan H.
====================
----------------------------